# Calibrating the SIDM Gravothermal Catastrophe with N-body Simulations

### v1

@@ I Introduction
Self-interacting dark matter (SIDM) is a class of dark matter models that include interactions between dark matter particles [1, 2, 3, 4, 5]. These models, which have theoretical motivation from particle physics [6, 7, 8, 9, 10, 11], may provide solutions to several unresolved contradictions between cold dark matter (CDM) models and observational data, such as the diversity problem (CDM; [12, 13, 14, 15, 16, 17, 18, 19], SIDM; [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]) and the cuspy halo problem (CDM; [31, 32, 33, 34, 35], SIDM; [36, 37]). There are many models for these self-interactions, some with elastic scattering and some with inelastic scattering. The inelastic models produce their own distinct phenomenology [38, 39, 40], but for this work we will focus solely on the elastic scattering case. In elastic models, the additional heat transfer channel provided by self-interactions causes the SIDM halos to undergo a time evolution not experienced by CDM halos, where an initially cuspy halo thermalizes and forms an isothermal core with a low, constant density [41]. The core formation process of an SIDM halo is followed by a slow increase in the halo’s core density, a phase called core-collapse. During this phase, the core transfers energy to the outer regions of the halo, causing particles to fall into the core. This collapse, also called the gravothermal catastrophe, is a runaway process that eventually causes a rapid density increase in the center of the halo [42, 43]. The phenomenon was first discovered in the study of globular clusters, where gravitational two-body scattering of stars provides the heat transfer instead of dark matter self-interaction [44], but the gravothermal catastrophe model has found success in describing dark matter halos as well.
@@ II.2 $\beta$ Prediction
$\begin{split}\kappa_{\mathrm{smfp}}&=\frac{2.1k_{B}v_{\mathrm{rms}}}{\sigma}\,,\\ \kappa_{\mathrm{lmfp}}&=0.27\beta nv_{rms}^{3}\sigma k_{B}/(Gm),\\ \end{split}$ (9)
@@
$\begin{split}\hat{\kappa}_{\mathrm{smfp}}&=\frac{2.1\hat{v}_{\mathrm{rms}}}{\widehat{\sigma}}\,,\\ \hat{\kappa}_{\mathrm{lmfp}}&=0.27\times 4\pi\beta\hat{\rho}\hat{v}_{\mathrm{rms}}^{3}\hat{\sigma}\,.\\ \end{split}$ (10)
@@ II.4 Measuring $\beta$
where A, C, D, and E are constants calibrated to a gravothermal fluid solution in the lmfp regime, and $\hat{t}=t\sqrt{4\pi G\rho_{s}}$. We use $A=-0.1078$, $C=-21.64$, $D=21.11$, and $E=2.238$ [51]. This model is valid from core formation until collapse, when $2.2381.341<\log_{10}(\beta_{\mathrm{eff}}\hat{\sigma}\hat{t})<$ 2.238, and we consider only this region of data in our fitting.
@@ III Results
Following Section II.4, the best fit $\beta_{\mathrm{eff}}$ for each individual simulation is found by fitting the central density (solid lines, $\hat{\sigma}$ designated by color) to Equation 23 (dashed black line). The only effect of changing $\beta_{\mathrm{eff}}$ in this figure is scaling the horizontal axis for an individual simulation, which will align that simulation to the gravothermal solution at the correct $\beta_{\mathrm{eff}}$. For this fit we only consider data after core formation and before critical core collapse, by the criterion $2.2381.341<\log_{10}(\beta_{\mathrm{eff}}\hat{\sigma}\hat{t})<$ 2.238. This is the effective range of Equation 23, and can be seen as the shaded region in Figure 1.
@@ Acknowledgements
This work was supported in part by a grant from the OSU College of Arts and Sciences Division of Natural and Mathematical Sciences. This work was also supported by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). Z.C. Zeng is partially supported by the Presidential Fellowship of the Ohio State University Graduate School.

### v2

@@ I Introduction
Self-interacting dark matter (SIDM) is a class of dark matter models that include interactions between dark matter particles [1, 2, 3, 4, 5]. These models, which have theoretical motivation from particle physics [6, 7, 8, 9, 10, 11], may provide solutions to several unresolved contradictions between cold dark matter (CDM) models and observational data, such as the diversity problem (CDM; [12, 13, 14, 15, 16, 17, 18, 19], SIDM; [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]) and the cuspy halo problem (CDM; [31, 32, 33, 34, 35, 36], SIDM; [37, 38]). There are many models for these self-interactions, some with elastic scattering and some with inelastic scattering. The inelastic models produce their own distinct phenomenology [39, 40, 41], but for this work we will focus solely on the elastic scattering case. In elastic models, the additional heat transfer channel provided by self-interactions causes the SIDM halos to undergo a time evolution not experienced by CDM halos, where an initially cuspy halo thermalizes and forms an isothermal core with a low, constant density [42]. The core formation process of an SIDM halo is followed by a slow increase in the halo’s core density, a phase called core-collapse. During this phase, the core transfers energy to the outer regions of the halo, causing particles to fall into the core. This collapse, also called the gravothermal catastrophe, is a runaway process that eventually causes a rapid density increase in the center of the halo [43, 44]. The phenomenon was first discovered in the study of globular clusters, where gravitational two-body scattering of stars provides the heat transfer instead of dark matter self-interaction [45], but the gravothermal catastrophe model has found success in describing dark matter halos as well.
@@ II.2 $\beta$ Prediction
[…]
@@
[…]
@@ II.4 Measuring $\beta$
where A, C, D, and E are constants calibrated to a gravothermal fluid solution in the lmfp regime, and $\hat{t}=t\sqrt{4\pi G\rho_{s}}$. We use A=0.1078A=-0.1078, C=21.64C=-21.64, $D=21.11$, and $E=2.238$ [52]. This model is valid from core formation until collapse, when $1.341<\log_{10}(\beta_{\mathrm{eff}}\hat{\sigma}\hat{t})<2.238$, and we consider only this region of data in our fitting.
@@ III Results
Following Section II.4, the best fit $\beta_{\mathrm{eff}}$ for each individual simulation is found by fitting the central density (solid lines, $\hat{\sigma}$ designated by color) to Equation 23 (dashed black line). The only effect of changing $\beta_{\mathrm{eff}}$ in this figure is scaling the horizontal axis for an individual simulation, which will align that simulation to the gravothermal solution at the correct $\beta_{\mathrm{eff}}$. For this fit we only consider data after core formation and before critical core collapse, by the criterion $1.341<\log_{10}(\beta_{\mathrm{eff}}\hat{\sigma}\hat{t})<2.238$. This is the effective range of Equation 23, and can be seen as the shaded region in Figure 1.
@@ Acknowledgements
This work was supported in part by a grant from the OSU College of Arts and Sciences Division of Natural and Mathematical Sciences. This work was also supported by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). S. Yang acknowledges support from the Director’s postdoctoral fellowship funded by the Laboratory Directed Research and Development (LDRD) program of Los Alamos National Laboratory (LANL) under project No. 20240863PRD2. Z.C. Zeng is partially supported by the Presidential Fellowship of the Ohio State University Graduate School.